Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b
(3,6); 8x+y=7

Can anyone help me understand this, I need help. Thank you so much

Answer by stanbon(75887)

Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b
(3,6); 8x+y=7
-----
Find the slope of the given line:
8x+y = 7
y = -8x+7
slope = -8
-------
ANY line perpendicular to the given line
must have a slope = 1/8.
------
You want the equation of a line with
slope = 1/8 passing thru (3,6)
-----
Form: y = mx + b
Substitute for y,x,and m and solve for "b":
6 = (1/8)3 + b
6 = 3/8 + b
(48/8) = (3/8) + b
b = 45/8
-----------
Equation:
y = (1/8)x + (45/8)
==========================
cheers,
Stan H.